OpenStudy (anonymous):

- 4x + 4/ 2y- 14 = -2/5

OpenStudy (solomonzelman):

Is it ( - 4x + 4)/( 2y- 14) = -2/5 or - 4x + 4/( 2y- 14) = -2/5 or (- 4x + 4)/( 2y)- 14 = -2/5 or - 4x + 4/ 2y- 14 = -2/5 Do you see the difference, you wrote it kind of unclear, which of those is it?

OpenStudy (anonymous):

d

OpenStudy (solomonzelman):

So it - 4x + 4/ 2y- 14 = -2/5

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

What do you need to do?

OpenStudy (solomonzelman):

put it in y=mx+b form?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

my answer is y= 5x + 2..

OpenStudy (anonymous):

is it correct? or did i miss a sign?

OpenStudy (solomonzelman):